### v1

# Shuffled Linear Regression via Spectral Matching

Abstract
Shuffled linear regression (SLR) seeks to estimate latent features through a linear transformation, complicated by unknown permutations in the measurement dimensions. This problem extends traditional least-squares (LS) and Least Absolute Shrinkage and Selection Operator (LASSO) approaches by jointly estimating the permutation, resulting in shuffled LS and shuffled LASSO formulations. Existing methods, constrained by the combinatorial complexity of permutation recovery, often address small-scale cases with limited measurements. In contrast, we focus on large-scale SLR, particularly suited for environments with abundant measurement samples. We propose a spectral matching method that efficiently resolves permutations by aligning spectral components of the measurement and feature covariances. Rigorous theoretical analyses demonstrate that our method achieves accurate estimates in both shuffled LS and shuffled LASSO settings, given a sufficient number of samples. Furthermore, we extend our approach to address simultaneous pose and correspondence estimation in image registration tasks. Experiments on synthetic datasets and real-world image registration scenarios show that our method outperforms existing algorithms in both estimation accuracy and registration performance.
Index Terms:
Shuffled linear regression, unlabelled sensing, image registration, correspondence estimation, permutation recovery.
I Introduction
Classical linear regression aims to recover multivariate signals from a noisy linear transformation, traditionally assuming that the correspondence between hidden signals and observations is perfectly known a priori. However, recent research has indicated that in many practical scenarios the observations are subject to an unknown permutation, significantly complicating the traditional linear regression formulation [SLR_ref3]. The challenge of permutation recovery has spurred the development of several solutions for the so-called shuffled linear regression (SLR) [SLR_ref3], also referred to as unlabeled sensing [8301561]. SLR aims to simultaneously recover hidden features and the associated permutations of measurements from the following model:
${\bf y}_{i}$ $=\boldsymbol{\Pi}^{\star}{\bf A}{\bf x}_{i}^{\star}+{\bf n}_{i},i=1,\cdots,t,$ (1)
where ${\bf y}_{i}\in\mathbb{R}^{m}$ and ${\bf x}_{i}^{\star}\in\mathbb{R}^{n}$ are the $i$-th measurement vector and latent feature vector, respectively, ${\bf A}\in\mathbb{R}^{m\times n}$ is a known sensing matrix, $\boldsymbol{\Pi}^{\star}\in\mathcal{P}_{m}$ is an unknown permutation matrix from the set of $m\times m$ permutation matrices $\mathcal{P}_{m}$, and ${\bf n}_{i}$ is an additive measurement noise vector independent of ${\bf x}_{i}$.
The model in (1) finds widespread use across various real-world applications. In data de-anonymization, attackers attempt to uncover hidden features observed through a linear transformation, intertwined with unknown measurement shuffling that a publisher performed before making the data public [SLR_ref5]. In image processing, SLR parallels the challenges encountered in classical image registration problems [10.1145/146370.146374], which entail pose and correspondence estimation [softposit]. This process involves aligning multiple point sets that capture the same objects at different snapshots. The aim of image registration is to identify the rigid transformation (referred to as the pose) and the permutation (referred to as the correspondence) between these datasets [Revisited]. Another significant application domain is the linear inversion system, particularly under conditions of signal asynchronization. For instance, signal sampling with jitter [1057717] illustrates a scenario where unknown observation times lead to unpredictably permuted measurements. This concern is particularly relevant in multi-target tracking and localization, which frequently involves estimating user positions from signals collected by asynchronous sensors [POORE20061074]. Similar challenges arise in molecular channels [7248460], where receptors receive molecular tokens at varied intervals, each bearing indistinguishable signatures.
Mathematically, for hidden features spanning real spaces as outlined in (1), the tasks of simultaneous signal recovery and permutation estimation can be approached through a variant of the least-squares (LS) method, often termed shuffled LS [SLR_ref3]. Furthermore, exploiting the structural properties of the hidden features can enhance the estimation accuracy. Notably, the sparsity of the input signals proves beneficial in solving under-determined linear inversion problems within the compressed sensing framework. A prominent method in SLR involves adapting LASSO algorithms to recover sparse signals while accommodating permutation estimation [SLR_CS3]. In the subsequent section, we will review related work addressing these aspects.
I-A Related Work
The SLR problem was initially explored in [7447086], where the authors established a necessary condition for exact permutation recovery in noiseless settings. The statistical limit for precise permutation recovery using an ideal maximum-likelihood estimator in a single-measurement scenario was subsequently investigated in [SLR_ref3]. In a related context, Ref. [SLR_ref7] analyzed the exact recovery conditions for shuffled LS algorithms with Gaussian sensing matrices. A branch-and-bound algorithm to solve shuffled LS was introduced in [pmlr-v97-tsakiris19a] and later enhanced by concave minimization techniques in [SLR_ref6]. Additionally, first-order methods aimed at enhancing computational efficiency were discussed in [SLR_ref4b, SLR_ref8]. Research in more general multiple-measurement scenarios, corresponding to $t>1$ in (1), was pursued in [SLR_ref2] for maximum-likelihood estimation. Furthermore, Ref. [pmlr-v119-zhang20n] built theoretical bounds for permutation estimation error under Gaussian sensing matrices and a restricted number of permutations.
In the context of sparse signal recovery, the authors in [SLR_CS3] developed a branch-and-bound algorithm for solving shuffled LASSO. To reduce the computational complexity, a gradient-based approach for estimating multiple-measurement vectors and unknown permutations was introduced in [SLR_CS2]. Additionally, Ref. [Alternating_Mat_SLR] proposed a permutation estimation algorithm based on spatial correlation matching and examined the impact of sparsity on estimation accuracy.
In addition to the standard SLR frameworks, various specialized variants have been developed to tackle different specific scenarios. A special case known as unlabeled de-noising, where ${\bf A}$ is set to an identity matrix in (1), was explored in [SLR_ref5]. An alternative strategy focusing on unlabeled sub-sampling was analyzed in [8234626]. Moreover, Ref. [SLR_ref1] introduced a shuffled total LS method tailored for situations with noisy observations of the sensing matrix. Additionally, research in [10.1214/18-EJS1498, SLR_SPermu_1] studied SLR problems involving sparse permutations, particularly under the assumption that only a small number of unknown permutations are possible in $\boldsymbol{\Pi}^{\star}$.
Image registration, a common application of SLR, is a fundamental problem in image processing. The challenges of estimating pose (rigid transformation) and correspondence (permutation) have been extensively studied, but typically in isolation, each with a long-established history. Notably, LS methods for pose estimation were discussed in-depth in [tpami, eggert1997estimating], while spectral matching techniques for correspondence estimation were explored in [1544893]. The combined challenge of simultaneous pose and correspondence estimation was first tackled in [softposit]. Since this initial exploration, various algorithms have been developed to address the joint estimation problems, including branch-and-bound methods [Revisited], graduated non-convexity methods [FGR], and semi-definite relaxations [9286491], among others.
I-B Our Contributions
The aforementioned work primarily focuses on scenarios with single measurements, corresponding to $t=1$ in (1), or few-shot measurements, where $t$ is comparable to other parameters such as $m$. Typically, these methods search for the permutation that aligns observations and hidden features within the spatial domain. In contrast, the isolated permutation recovery challenges have been extensively addressed through spectral matching approaches that aim to align the eigenbases of the input and output data [GMP_eign1]. These methods are acknowledged for their computational efficiency and robustness against noise, particularly when a substantial number of measurement samples are available [GMP_eign3].
Motivated by these considerations, in this work, we explore spectral matching to solve the SLR problems under the large-system limit. We consider the setup where the number of measurements $t$ significantly exceeds the dimension of measurements $m$. Our approach utilizes measurement samples to estimate the eigenbases of the signal sample covariance matrix. We then determine the optimal permutation by aligning the spectral components through spectral matching. This alignment allows us to reformulate the problem into traditional least squares (LS) or LASSO optimization frameworks. Finally, we extend our method to address the practical challenge of three-dimensional (3D) image registration. The key contributions of our work are summarized as follows:
  • $\bullet$
    We propose spectral matching methods for shuffled LS and shuffled LASSO problems and demonstrating a reduced computational complexity compared to existing algorithms.
  • $\bullet$
    We show that the permutation estimation error of our proposed method diminishes at the rate of $\mathcal{O}(1/\sqrt{t})$ when $t\gg m^{3}$. By comparing to the state-of-the-art algorithm [pmlr-v119-zhang20n], which computes the permutation in the spatial domain, our analysis underscores the superior performance of spectral matching, particularly when a large number of observation samples are available.
  • $\bullet$
    Based on permutation error analysis, we study the estimation error for both shuffled LS and LASSO. Our method achieves sub-linear error rates, balancing efficiency with optimal asymptotic order as $t\to\infty$. Specifically, as $t\to\infty$, the estimation error of our method achieves the optimal asymptotic order, albeit with a slower convergence rate ($\mathcal{O}(1/\sqrt{t})$) compared to the optimal but computationally prohibitive solver whose rate is $\mathcal{O}(1/t)$.
  • $\bullet$
    We adapt the proposed method for 3D image registration by combining spectral matching with a translation estimation step. While image registration does not typically follow the large-system limit, the orthogonality of the rotation matrix enables accurate covariance estimation, broadening our method’s applicability.
We conduct experiments on synthetic data and real-world 3D point datasets for addressing both shuffled LS and shuffled LASSO problems. The results demonstrate that our proposed method surpasses state-of-the-art baselines in performance and achieves accurate permutation estimation, particularly when the sample size $t$ is sufficiently large.
I-C Organization and Notations
The remainder of this paper is organized as follows. We formulate the SLR problems in Section II. In Section III, we develop the spectral matching algorithm. In Section IV, we analyze the estimation error of the proposed method. In Section LABEL:sec5, we adapt the proposed method to solve 3D image registration problems. In Section LABEL:sec6, we present numerical results to evaluate the proposed method. Finally, this paper concludes in Section LABEL:sec7.
Throughout, we use regular, bold small, and bold capital letters to denote scalars, vectors, and matrices, respectively. We use ${\bf X}^{T}$ to denote the transpose of matrix ${\bf X}$, $\mathop{\rm tr}\nolimits({\bf X})$ to denote its trace, and $\mathop{\rm rank}\nolimits({\bf X})$ to denote its rank. We use $x_{i}$ to denote the $i$-th entry of vector ${\bf x}$, $x_{ij}$ or $[{\bf X}]_{i,j}$ interchangeably to denote the $(i,j)$-th entry of matrix ${\bf X}$, and ${\bf x}_{j}$ to denote the $j$-th column of ${\bf X}$. The real normal distribution with mean $\boldsymbol{\mu}$ and covariance ${\bf C}$ is denoted by $\mathcal{N}(\boldsymbol{\mu},{\bf C})$, and the cardinality of set $\mathcal{S}$ is denoted by $\lvert\mathcal{S}\rvert$. We use $\lVert\cdot\rVert_{p}$ to denote the $\ell_{p}$ norm, $\lVert\cdot\rVert_{F}$ (resp. $\lVert\cdot\rVert_{2}$) to denote the Frobenius (resp. spectral) matrix norm, ${\bf I}_{n}$ to denote the $n\times n$ identity matrix, $\bf 1$ (resp. $\bf 0$) to denote the all-one (resp. all-zero) vector with an appropriate size, and $\mathop{\rm diag}\nolimits({\bf x})$ to denote a diagonal matrix with the diagonal entries specified by ${\bf x}$. For two scalars $a$ and $b$, we denote $a\gtrsim b$ or $b\lesssim a$ interchangeably if there exists a positive constant $c_{0}$ such that $a\geq c_{0}b$. We denote $a\asymp b$ when both $a\gtrsim b$ and $b\gtrsim a$ hold.
II Problem Formulation
We study the SLR model described in (1). Our goal is to estimate both the permutation $\boldsymbol{\Pi}^{\star}$ and the latent features ${\bf X}^{\star}=[{\bf x}_{1}^{\star},\cdots,{\bf x}_{t}^{\star}]$ from the noisy measurements ${\bf Y}=[{\bf y}_{1},{\bf y}_{2},\cdots,{\bf y}_{t}]$ and the sensing matrix ${\bf A}$. Unless otherwise specified, the feature vectors $\{{\bf x}_{i}^{\star}\}_{i=1}^{t}$ and the noise vectors $\{{\bf n}_{i}\}_{i=1}^{t}$ are modeled as unknown independent and identically distributed (i.i.d.) random vectors drawn from unspecified distributions. These vectors are assumed to have zero means with given positive semidefinite covariance matrices $\mathbb{E}[{\bf x}_{i}^{\star}({\bf x}_{i}^{\star})^{T}]={\bf C}_{X}$ and $\mathbb{E}[{\bf n}_{i}{\bf n}_{i}^{T}]={\bf C}_{N}$. The adaptation of the proposed estimation approach to accommodate features with non-zero means and certain dependent samples will be explored in Section LABEL:sec5.
Under the condition of $m\geq n$, the SLR estimation problem has been explored through a shuffled LS formulation [SLR_ref6]:
$(\widehat{\boldsymbol{\Pi}},\widehat{\bf X})\in\operatorname*{\mathop{\arg\min}}_{{\bf X}\in\mathbb{R}^{d\times t},\boldsymbol{\Pi}\in\mathcal{P}_{m}}\lVert\boldsymbol{\Pi}^{T}{\bf Y}-{\bf A}{\bf X}\rVert_{F}^{2}.$ (2)
It has been shown that the problem in (2) is NP-hard [SLR_ref3].
In addition to the shuffled LS, we examine a variant of this problem that accommodates underdetermined linear regression with $n\geq m$, specifically when the input features in ${\bf X}$ are sparse. This setup generalizes the well-established compressed sensing framework [1614066] with the additional complexity of an unknown measurement permutation. The sparsity assumption models scenarios where signals are sparsely represented under certain linear transformations, such as image pixels under wavelet or Fourier transforms. Previous work [SLR_CS3, Alternating_Mat_SLR] proposed to solve compressed sensing under unknown permutations using the following $\ell_{1}$-regularized formulation:
$(\widehat{\boldsymbol{\Pi}},\widehat{\bf X})\in\operatorname*{\mathop{\arg\min}}_{{\bf X}\in\mathbb{R}^{d\times t},\boldsymbol{\Pi}\in\mathcal{P}_{m}}\lVert\boldsymbol{\Pi}^{T}{\bf Y}-{\bf A}{\bf X}\rVert_{F}^{2}+\rho\lVert{\bf X}\rVert_{1,1},$ (3)
where $\rho$ is a tunable penalty coefficient and $\lVert{\bf X}\rVert_{1,1}=\sum_{i,j}x_{ij}$ denotes the entrywise $\ell_{1}$-norm. We refer to (3) as shuffled LASSO regression.
The goal of this work is to design computationally efficient algorithms for solving (2) and (3), for cases where a substantial number of measurement samples is available, i.e., $t\gg m$.
Remark 1.
We note that the problem in (3) is fundamentally different from those considered in [10.1214/18-EJS1498, SLR_SPermu_1]. The latter studies typically involve a dense feature matrix ${\bf X}$ and a sparse permutation $\boldsymbol{\Pi}^{\star}$. In contrast, our formulation does not impose any restrictions on the unknown permutation. Consequently, in the worst-case scenario, there could be $m!$ possible random permutations to search.
III Spectral Matching
In this section, we present an approach, with relatively low-complexity, to solve (2) and (3), which is particularly suitable for the high-dimensional scenarios with $t\gg m$. The main challenge in these combinatorial problems stems from the unknown permutation matrix, casting the issue as a quadratic assignment problem, as discussed in [SLR_CS2]. Once a solution for the permutation $\widehat{\boldsymbol{\Pi}}$ is identified, the solution to ${\bf X}$ in (2) exhibits a closed-form expression as
$\widehat{\bf X}(\widehat{\boldsymbol{\Pi}})={\bf A}^{\dagger}\widehat{\boldsymbol{\Pi}}^{T}{\bf Y},$ (4)
where ${\bf A}^{\dagger}=({\bf A}^{T}{\bf A})^{-1}{\bf A}^{T}$ is the pseudo-inverse of ${\bf A}$. With the LASSO formulation, once $\widehat{\boldsymbol{\Pi}}$ is determined, problem (3) can be mapped into:
$\widehat{\bf X}(\widehat{\boldsymbol{\Pi}})=\operatorname*{\mathop{\arg\min}}_{{\bf X}\in\mathbb{R}^{d\times t}}\lVert\widehat{\boldsymbol{\Pi}}^{T}{\bf Y}-{\bf A}{\bf X}\rVert F^{2}+\rho\lVert{\bf X}\rVert_{1,1},$ (5)
which is a convex programming problem and can be efficiently solved using off-the-shelf optimization solvers.
In the subsequent analysis, we focus on solving for $\widehat{\boldsymbol{\Pi}}$ from the model in (1). Our approach hinges on estimating the covariance matrix of ${\bf Y}$ and aligning its eigenvectors with those of ${\bf C}_{X}$. Specifically, we denote the covariance of ${\bf y}_{i}$ by ${\bf C}_{Y}=\mathbb{E}[{\bf y}_{i}{\bf y}_{i}^{T}]$. From (1), the covariance is expressed as:
${\bf C}_{Y}=\boldsymbol{\Pi}^{\star}{\bf A}{\bf C}_{X}{\bf A}^{T}(\boldsymbol{\Pi}^{\star})^{T}+{\bf C}_{N}.$ (6)
Consequently, ${\bf C}_{Y}-{\bf C}_{N}$ equals to ${\bf A}{\bf C}_{X}{\bf A}^{T}$ subject to the permutation of $\boldsymbol{\Pi}^{\star}$ applied to its rows and columns. We detail their eigendecompositions as:
${\bf C}_{Y}-{\bf C}_{N}$ $=(\boldsymbol{\Pi}^{\star}{\bf V})\boldsymbol{\Lambda}(\boldsymbol{\Pi}^{\star}{\bf V})^{T},$
${\bf A}{\bf C}_{X}{\bf A}^{T}$ $={\bf V}\boldsymbol{\Lambda}{\bf V}^{T},$ (7)
where $\boldsymbol{\Lambda}=\mathop{\rm diag}\nolimits(\lambda_{1},\cdots,\lambda_{m})$ is a diagonal matrix containing the eigenvalues in descending order, and ${\bf V}$ is an orthogonal eigenmatrix of ${\bf A}{\bf C}_{X}{\bf A}^{T}$.
As a result, the eigenvectors of ${\bf C}_{Y}-{\bf C}_{N}$ correspond to those of ${\bf A}{\bf C}_{X}{\bf A}^{T}$, subject to the permutation $\boldsymbol{\Pi}^{\star}$. Therefore, $\boldsymbol{\Pi}^{\star}$ can be determined by aligning their eigenvector matrices. To facilitate this, we compute the sample covariance matrix of ${\bf Y}$, denoted by $\widehat{\bf C}_{Y}$, to approximate ${\bf C}_{Y}$ and represent the corresponding eigendecomposition as:
$\widehat{\bf C}_{Y}-{\bf C}_{N}=\frac{1}{t}\sum_{i=1}^{t}{\bf y}_{i}{\bf y}_{i}^{T}-{\bf C}_{N}={\bf U}{\bf S}\widehat{\boldsymbol{\Lambda}}{\bf S}{\bf U}^{T},$ (8)
where $\widehat{\boldsymbol{\Lambda}}=\mathop{\rm diag}\nolimits(\hat{\lambda}_{1},\cdots,\hat{\lambda}_{m})$ consists of the sample eigenvalues in descending order, ${\bf U}$ is the sample eigenvector matrix, and ${\bf S}$ is a diagonal matrix with entries of either $1$ or $-1$. The matrix ${\bf S}$ accounts for the inherent sign ambiguity in eigendecompositions. [^1] If $({\bf u},\lambda)$ is an eigenpair for a matrix, then $(-{\bf u},\lambda)$ is also a valid eigenpair for the same matrix.
With a large $t$, the sample covariance $\widehat{\bf C}_{Y}$ provides a good approximation of ${\bf C}_{Y}$. By combining (7) and (8), we see that ${\bf U}$ closely approximates $\boldsymbol{\Pi}^{\star}{\bf V}{\bf S}$. This motivates us to compute the permutation by minimizing the distance between ${\bf U}$ and $\boldsymbol{\Pi}^{\star}{\bf V}{\bf S}$. To tackle the sign ambiguity, we compute the permutation by solving the following problem (cf. [GMP_eign1, BlindGM]):
$\widehat{\boldsymbol{\Pi}}=\operatorname*{\mathop{\arg\max}}_{\boldsymbol{\Pi}\in\mathcal{P}_{m}}\mathop{\rm tr}\nolimits\left(abs\left({\bf V}_{\mathcal{A}}\right)abs\left({\bf U}^{T}_{\mathcal{A}}\right)\boldsymbol{\Pi}\right),$ (9)
where ${abs}(\cdot)$ is the matrix containing the absolute values of the entries in the input matrix, and ${\bf V}_{\mathcal{A}}$ denotes the submatrix consisting of the columns of ${\bf V}$ indexed by a subset $\mathcal{A}\subset\{1,2,\cdots,m\}$. Problem (9) is a linear assignment problem and can be efficiently solved using the Hungarian method [Hungarian].
To ensure precise alignment of eigenvectors, we propose to select $\mathcal{A}$ from the principal components of ${\bf V}$ that correspond to distinct and non-zero eigenvalues as
$\mathcal{A}\subset\mathcal{C}\triangleq\{1\leq i\leq m:\lambda_{i}>0,\lambda_{i}\neq\lambda_{j},\forall j\neq i\}.$ (10)
Given that $\mathop{\rm rank}\nolimits({\bf A}{\bf C}_{X}{\bf A}^{T})\leq\min\{m,n\}$, we have $|\mathcal{A}|\leq|\mathcal{C}|\leq\min\{m,n\}$. Although one could set $|\mathcal{A}|=|\mathcal{C}|$, it has been shown in [BlindGM] that selecting only a subset of principal components, particularly those corresponding to large spectral gaps, boosts estimation accuracy. For instance, $\mathcal{A}$ can be progressively computed using the line-search algorithm in [BlindGM, Algorithm 3]. The selection of $\mathcal{A}$ from the principal components guarantees a strictly positive spectral gap among the chosen spectral components. Specifically, for any given $\mathcal{A}$, the minimum spectral gap is denoted by:
$\delta\triangleq\min_{i\in\mathcal{A}}\{\lambda_{i}-\lambda_{i+1}\},$ (11)
where $\lambda_{0}=\infty$ and $\lambda_{m+1}=0$ are defined for notational convenience. The uniqueness of the selected eigenvalues ensures that $\delta>0$. This condition is crucial for ensuring the correct alignment of the sample eigenvectors, which is a prerequisite for achieving accurate permutation estimation.
In summary, our approach first computes the sample covariance matrix and its eigendecomposition by (7) and (8). Subsequently, we estimate $\widehat{\boldsymbol{\Pi}}$ by solving the linear assignment problem in (9). The solution for ${\bf X}$ is then derived either directly from (4) or by solving the resultant convex problem in (5). The total computational complexity for computing (7)–(9) scales at the rate of $\mathcal{O}(m^{2}(m+t))$. In contrast, the state-of-the-art approximate branch-and-bound algorithm in [SLR_ref6], applicable only when $t=1$, exhibits a prohibitive complexity of $\mathcal{O}(n^{2m})$ for large values of $n$. Additionally, the complexity of the spatial-domain matching algorithm in [pmlr-v119-zhang20n] grows at a rate of $\mathcal{O}(m^{2}(m+n+t))$.
IV Performance Analysis
We analyze the estimation error for the proposed solutions to shuffled LS and shuffled LASSO. As detailed in Section II, we focus on the regime where $t\gg m$.
IV-A Permutation Estimation Error Analysis
We begin by examining the permutation estimation error for the solution in (9). Denote $k\triangleq|\mathcal{A}|\leq\min\{m,n\}$. To assess the accuracy of the estimate $\widehat{\boldsymbol{\Pi}}$, we introduce a noiseless permutation estimation problem as a counterpart to (9):
$\max_{\boldsymbol{\Pi}\in\mathcal{P}_{m}}\mathop{\rm tr}\nolimits\left(\boldsymbol{\Pi}^{T}\boldsymbol{\Pi}^{\star}{\bf V}_{\mathcal{A}}{\bf V}^{T}_{\mathcal{A}}\right),$ (12)
where the sign ambiguity no longer exists, thus negating the need for the absolute value operation. Problem (12) mirrors the ideal scenario where the covariance estimation in (8) is perfectly noiseless, i.e., $\widehat{\bf C}_{Y}={\bf C}_{Y}$. We note that (12) is an oracle formulation as it presupposes the knowledge of $\boldsymbol{\Pi}^{\star}$. It can be verified that for any $\boldsymbol{\Pi}\in\mathcal{P}_{m}$,
$\mathop{\rm tr}\nolimits\left(\boldsymbol{\Pi}^{T}\boldsymbol{\Pi}^{\star}{\bf V}_{\mathcal{A}}{\bf V}^{T}_{\mathcal{A}}\right)\leq\mathop{\rm tr}\nolimits\left({\bf V}_{\mathcal{A}}{\bf V}^{T}_{\mathcal{A}}\right)=k,$ (13)
where equality holds if $\boldsymbol{\Pi}=\boldsymbol{\Pi}^{\star}$. In other words, the true permutation $\boldsymbol{\Pi}^{\star}$ maximizes the objective in (12). Hence a measure of the permutation accuracy of our solution $\widehat{\boldsymbol{\Pi}}$ is the normalized optimality gap with respect to (12), defined as
$1-\frac{1}{k}\mathop{\rm tr}\nolimits\left(\widehat{\boldsymbol{\Pi}}^{T}\boldsymbol{\Pi}^{\star}{\bf V}_{\mathcal{A}}{\bf V}^{T}_{\mathcal{A}}\right)\in[0,1].$ (14)
Intuitively, (14) quantifies the average mismatched ratio for the proposed solution, which equals to zero if $\widehat{\boldsymbol{\Pi}}=\boldsymbol{\Pi}^{\star}$. The following theorem bounds this optimality gap.